formula 2x+2y The production possibilities curve for an economy is given by the 150, where x and y are quantities of produced goods, while the tility function in this economy is U(x.y) xy The utility level in a competitive equilibrium will be:
Answer #1

At the tangency of utility function and PPF, slopes are equal

2 * 2x / 2 * 2y = y / x

x^2 = y^2

Use this in PPF

2y^2 + 2y^2 = 150

y = (150/4)^0.5 = 6.1237

This also gives x = 6.1237

Hence the utility level is U = 6.1237*6.1237 = 37.5.